2ln(100-3x)-8=0.5
Round answer to nearest thousandth,solve for x.

Answer by stanbon(75887)

2ln(100-3x)-8=0.5
Divide thru by 2 to get::
ln(100-3x) -4 = 0.25
ln(100-3x) = 4.25
----
100-3x = e^4.25
3x = 100-e^4.25
3x = 29.895
x = 9.965
-----
Round answer to nearest thousandth,solve for x.
---------------
Cheers,
Stan H.
--------------